Solve by elimination method:
5x+5y= -13
7x-3y= 17

Found 2 solutions by sowmya, aswathytony:
Answer by sowmya(32)

Solve by elimination method:
5x+5y= -13
7x-3y= 17

eq 1:
x+y = -13/5
7x = 17+3y
x = 17/7 + 3/7 y
= 17/7 + 3/7 * (-x-13/7)
=17/7 -3x/7 -39/49
10x/7 = 119-39/49
= 80/49
x = 8/7
y = -131/35

Answer by aswathytony(47)

5x + 5y = -13 .......(1)
7x - 3y = 17 ............(2)

(1) * 3 & (2) * 5
5x * 3 + 5y * 3 = -13* 3
15 x + 15 y = - 39 ............(3)
7x * 5 - 3y * 5 = 17 *5
35 x - 15 y = 85 ..........(4)
(3) + (4)
50x = 46

x = 46 /50 = 0.92
substituting x =0.92 in (1)
5 * 0.92 + 5y = -13
4.6 + 5y = -13

5y = -13 - 4.6 = -17.6
y = -17.6 / 5 = -3.52